OpenStudy (anonymous):

If there is an inverse relationship between x and y, and the constant of variation is 42, what is y when x = 3?

OpenStudy (anonymous):

@theonlycool13

OpenStudy (anonymous):

he here to help :-0

OpenStudy (anonymous):

there are two possibilities... either.. x = k/y or y = k/x now when x = 3 and k = 42... then ==> 3 = 42/y ==> y = 14

OpenStudy (anonymous):

is this right? @AnswerMyQuestions

OpenStudy (anonymous):

I think. Thanks.

OpenStudy (anonymous):

u r welcom